OpenStudy (anonymous):

The line above goes through points (-4,-1)and(0,2). What is the equation of the line? A. Y=(3/4)x-2 B. Y=(3/2)x+3 C. Y=(4/3)x+2 D. Y=(3/4)x+2

OpenStudy (anonymous):

@garrett_payne think you could explain this to me as well?

OpenStudy (anonymous):

very similar thing here. find slope then intercept

OpenStudy (anonymous):

How would I do that for this one?

OpenStudy (anonymous):

pick point 1 and point 2 then find (Y2-y1)/(x2-X1) remember it's written as (X,Y)

OpenStudy (anonymous):

so -4 - -1 -------- 0 - 2

OpenStudy (anonymous):

??

OpenStudy (anonymous):

no you're writing it as (x1-y1) /(x2-y2)

OpenStudy (anonymous):

first write what Y1,Y2, X1 and X2 are

OpenStudy (anonymous):

Y1=-4,Y2=-1, X1=0,X2=2

OpenStudy (anonymous):

no, the numbers are read as (x,Y) so say (-4,-1) is point 1 then -4 is x and -1 is y

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

\[m = y _{2}-y _{1}/x _{2}-x _{1}\]

OpenStudy (anonymous):

where m is the gradient of the line, try doing it now :)

OpenStudy (anonymous):

how are you doing with this @kaek98

OpenStudy (anonymous):

Not great

OpenStudy (anonymous):

I dont understand any of this

OpenStudy (anonymous):

look at my last post call it point one then do the same for point 2

OpenStudy (anonymous):

so 0 is x and 2 is y

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

now that you know Y1, X1, Y2 and X2 plug them in the slope equation I showed you

OpenStudy (anonymous):

(-4-0)/(-1-2)

OpenStudy (anonymous):

close but backwards (-1-2)/(-4-0) remember x is first then y

OpenStudy (anonymous):

slope is -3/-4 or 0.75

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

that value is m (slope) remember y=mx+b plug in either point say (0,2) y=2 m=3/4 x=0 and find b

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Am I right? @garrett_payne

OpenStudy (anonymous):

yes you are. You can do it just try to think about what the numbers mean and not get flustered because you think you're bad a math

OpenStudy (anonymous):

Thanks :)